I am supposed to solve this problem by completing the square:
(x squared) + 11x + 10=0

Answer by LifeDeathnCandy(5)

x^2+ 11x + 10=0
To solve this you have to find the x coordinates. To do that you have to use the quadratic equation.( x=-b+-(the square root of)b^2-4ac over 2a)
To start of you have to solve for the b^2-4ac first.
11^2-4(1)(10)=81
So then you find the square root of 81 and that is 9.
So you do -11+9/2=-1
And -11-9/2=-10.
So your two x coordinates are (-1,0) and (-10,0)
Then you have to find the vertex using -b/2a. -11/2=-5.5
Vertex= -5.5 <----xcoordinate of the vertex
From here you substitute the x coordinate of the vertex into the original equation.
-5.5^2+11(-5.5)+10=30.25-60.5+10=-20.25<----ycoordinate of the vertex.
Vertex=(-5.5,-20.25)
So now you have your xcoordinates and your vertex and from here you can graph it if you have to.
Hope this helps.